# Sound waves from primordial black hole formations

### v1

@@ III.4 Energy of sound shells
Since we consider compensated perturbations and enforce outer boundary conditions matching the FLRW solution, the energy of the PBH must be balanced by the energy of the overdense and underdense shells, satisfying the relation $E_{+}+E_{-}\approx M_{\mathrm{PBH}}$. Although a high-density region exists near the PBH that does not propagate outwards, its contribution to the total energy is negligible due to its small radial extent, ensuring that this energy balance holds precisely. Furthermore, this relation is generally valid across various PBH formation scenarios, as discussed in Ref. Deng (2021). Consequently, a compensating underdense structure must always be present near PBHs, whereas the overdense structure arises only in near-critical cases.
IV Conclusions and discussion
In this work, we have numerically investigated PBH formation from super-horizon curvature perturbations and the associated generation and propagation of compression and sound waves in the early Universe. Using the Misner-Sharp formalism with an excision technique, we have explored a broad range of scenarios—from subcritical to supercritical perturbations—and examined the effects of varying equations of state and perturbation profiles.
Our results reveal that the subsequent evolution after PBH formation is highly sensitive to the perturbation amplitude. For amplitudes slightly above the critical threshold, a bounce occurs, generating a compression wave composed of both overdense and underdense shells. The rarefaction region between the PBH and the overdense shell effectively cuts off further accretion, leading to a smaller PBH mass in near-critical cases. In contrast, for significantly supercritical perturbations, only an underdense shell is observed to propagate outwards. Furthermore, we find that softer equations of state or steeper perturbation profiles inhibit the formation of compression waves, consistent with the expectation that reduced pressure gradients diminish the ability of pressure forces to overcome gravity.
As the sound shells propagate far from the PBH, they behave as linear perturbations on a flat FLRW background. The linearized conservation equations describe their propagation, demonstrating that the comoving thickness of the shells remains approximately constant—a result corroborated by our numerical simulations. Moreover, we have established an empirical relation between the comoving thickness of the shells and the Hubble radius at horizon re-entry. This relation provides a means to link the PBH mass with the peak frequency of the gravitational wave spectrum, which can be used to constrain PBH abundance in specific mass ranges.
Furthermore, we find that the energy contained within the overdense shell, measured via the Misner-Sharp mass, decreases as the perturbation amplitude departs from the critical threshold, following a scaling law analogous to that of the PBH mass but with an opposite prefactor and an offset. Additionally, the energy of the PBH and the overdense shell must be balanced by that of the underdense shell. As a result, a compensating underdense structure is always present near PBHs, while an overdense structure forms only in near-critical cases.
Although our numerical simulations provide valuable insights into the dynamics of PBH formation and sound wave evolution, further research is needed. In this study, we have considered only generalized Gaussian curvature perturbations. However, curvature perturbations derived from the primordial power spectrum can exhibit more complex features, such as multi-peak or oscillatory profiles, particularly in the presence of non-Gaussianities Franciolini et al. (2018); Atal and Germani (2019); Atal et al. (2019); Cai et al. (2018); Atal et al. (2020); Escrivà et al. (2022). These variations may significantly modify sound wave profiles, leading to different gravitational wave spectra—a topic that merits further investigation. Future studies should also examine non-spherical effects Yoo et al. (2020); Yoo (2024); Escrivà and Yoo (2024a, b) and alternative PBH formation mechanisms to assess the robustness of our conclusions across a wider range of cosmological scenarios.
Moreover, simulating the formation of multiple PBHs and the interactions between their associated sound waves would offer a critical test of our proposed gravitational wave generation mechanism and the applicability of the sound shell model. This scenario, reminiscent of the black hole lattice universe Yoo et al. (2012, 2013); Yoo and Okawa (2014); Bentivegna and Korzynski (2012); Korzyński (2014); Clifton et al. (2014); Bibi et al. (2017); Bentivegna et al. (2018); Yoo et al. (2019); Coley (2020), is left for future work.
In summary, our study provides a comprehensive analysis of PBH formation and the subsequent evolution of sound waves, shedding new light on the interplay between nonlinear gravitational collapse, fluid dynamics, and early Universe cosmology. These results enhance our understanding of PBH formation and pave the way for future research into their observable cosmological consequences.
Acknowledgements.
This work is supported by the National Key Research and Development Program of China Grant No. 2021YFC2203004, No. 2021YFA0718304, and No. 2020YFC2201501, the National Natural Science Foundation of China Grants No. 12422502, No. 12105344, No. 12235019, No. 12047503, No. 12073088, No. 11821505, No. 11991052, and No. 11947302, and the Science Research Grants from the China Manned Space Project with No. CMS-CSST-2021-B01 (supported by China Manned Space Program through its Space Application System). HD was supported by Yuri Levin’s Simons Investigator Grant PG012519.
Appendix A Convergence test
In this appendix, we present a convergence test for our numerical code. To quantify the numerical error, we use the $L_{2}$-norm of the Hamiltonian constraint equation (5b), which should vanish if the Einstein and conservation equations are solved correctly. Specifically, we define the Hamiltonian constraint violation and its $L_{2}$-norm as
$\mathcal{H}$ $\equiv\frac{D_{r}M}{4\pi\Gamma\rho R^{2}}-1,$ (56a)
${||\mathcal{H}||}_{2}$ $\equiv\frac{1}{N}\sqrt{\sum_{i=1}^{N}{|\mathcal{H}_{i}|}^{2}}.$ (56b)
In terms of our redefined variables, the Hamiltonian constraint violation is expressed as
$\mathcal{H}=\frac{\bar{M}+\frac{\bar{r}\bar{R}}{3(\bar{r}\bar{R})^{\prime}}\bar{M}^{\prime}}{\bar{\rho}}-1.$ (57)
To eliminate errors arising from the use of the long-wavelength approximation for the initial conditions, we subtract the initial Hamiltonian constraint violation at each time step.
We use the near-critical perturbation ($\delta_{m}=0.5$) as a test case. The code is run at various grid resolutions, with $N=1500,2000,2500,3000,3500,4000,4500,5000,5500,6000$, and the logarithmic grid constant is set to $C=1+9/N$ for all cases. All other parameters are identical to those used in the main text. The left panel of Fig. 12 shows the time evolution of the $L_{2}$-norm of the Hamiltonian constraint violation. It can be seen that, for sufficiently high grid resolution, the violation remains at approximately $10^{-3}$ throughout the simulation (extending much later than in previous studies). The right panel of Fig. 12 plots the $L_{2}$-norm at the end of the simulation as a function of the grid resolution. The black dashed line, representing a $N^{-4}$ scaling, fits the data well, confirming the fourth-order convergence of the central difference method employed.
Figure 12: Convergence test of the numerical code. Left panel: Time evolution of the $L_{2}$-norm of the Hamiltonian constraint violation for various grid resolutions. Right panel: $L_{2}$-norm of the Hamiltonian constraint violation at the end of the simulation as a function of the grid resolution. The black dashed line indicates the expected $N^{-4}$ scaling.

### v3

@@ III.4 Energy of sound shells
Here $\int_{r_{\mathrm{PBH}}}^{r_{A}}4\pi R^{2}(\rho-\rho_{b})\mathrm{d}R$ is the energy of the high-density region near the PBH that does not propagate outwards; its contribution to the total energy is negligible due to its small radial extent. Besides, we have $\Delta M(r_{\mathrm{PBH}})=M(r_{\mathrm{PBH}})-M_{b}(r_{\mathrm{PBH}})\approx M(r_{\mathrm{PBH}})=M_{\mathrm{PBH}}$, since the background mass within $r_{\mathrm{PBH}}$ is negligible compared to $M_{\mathrm{PBH}}$. These relations have been verified by our numerical results. Furthermore, this balance formula is generally valid across various PBH formation scenarios, as discussed in Ref. 2106.09817. Consequently, a compensating underdense structure must always be present near PBHs, whereas the overdense structure arises only in near-critical cases.
III.5 Effects of the curvature perturbation profile
It is well-known that the formation threshold 1310.3007; 1809.02127; 1907.13311; 2007.05564, mass 2103.03867, and abundance 1805.04087 of PBHs depend on the shape of the curvature perturbations. Therefore, it is of interest to study how the formation of compression waves depends on the perturbation profile. To this end, we follow Refs. 1809.02127; gr-qc/0605122 to consider two more general profiles. First, we examine a family of generalized Gaussian profiles:
$K(r)=\mathcal{A}\left(\frac{r}{\Delta}\right)^{2\lambda}\exp\left[-\frac{1}{2}\left(\frac{r}{\Delta}\right)^{2\beta}\right],$ (58)
where $\Delta$ is related to the perturbation scale $r_{i}$ by
$\Delta=\left(\frac{2(\lambda+1)}{\beta}\right)^{-1/2\beta}r_{i}.$ (59)
Here, $\beta$ and $\lambda$ are parameters that control the shape of the perturbation: $\beta$ determines the steepness of the profile, while $\lambda$ controls the location of the peak. Eq. (35) is recovered when $\beta=1$ and $\lambda=0$. We consider three cases: a steeper profile with $\beta=0.5,\lambda=0$; a flatter profile with $\beta=4,\lambda=0$; and an off-centered profile with $\beta=1,\lambda=4$. The curvature perturbation profiles for these three cases are shown as solid lines in Fig. 11. For comparison, the Gaussian profile is also shown as a dotted line. The corresponding formation thresholds are $\delta_{c}\approx 0.46695\pm 1\times 10^{-5}$ (steeper), $\delta_{c}\approx 0.56837\pm 1\times 10^{-5}$ (flatter), and $\delta_{c}\approx 0.59013\pm 1\times 10^{-5}$ (off-centered).
Figure 11: Curvature perturbation profiles for different cases: Gaussian profile (dotted line); generalized Gaussian profiles (solid lines) with $\beta=0.5,\lambda=0$ (blue), $\beta=4,\lambda=0$ (orange), and $\beta=1,\lambda=4$ (green); and smoothed top-hat profiles (dashed lines) with $\Delta_{1}=4R_{H}$ (red) and $\Delta_{1}=8R_{H}$ (purple). All profiles are normalized such that $\mathrm{max}(K)=1$.
In all cases, the perturbation amplitudes are chosen to satisfy $\delta_{m}-\delta_{c}\approx 0.00226$ as well. Fig. 12 shows the time evolution of the energy density contrast for these different profiles. The PBH forms at $t\approx 31.7t_{m}$, $t\approx 18.7t_{m}$, and $t\approx 19.8t_{m}$, respectively. The flatter profile leads to a larger pressure gradient and, therefore, a higher formation threshold than the steeper profile. As can be seen in the top panel of the figure, the flatter profile produces a stronger bounce (as indicated by the red lines), reflecting an enhanced pressure gradient, while the steeper profile results in a wider compression wave. For smaller or larger $\beta$, our code is unable to follow the dynamical evolution for extended periods, and we leave the detailed study of the dependence of the compression wave shape on the perturbation profile for future work. For the off-centered profile, simulations indicate that during the early evolution, matter redistributes to fill the central depression, converging toward a centrally peaked profile (see the bottom-left panel of the figure). Consequently, the resulting compression wave is similar to that of a centrally peaked profile (as shown in the bottom-right panel of the figure).
Figure 12: Time evolution of the energy density contrast for generalized Gaussian profiles: steeper profile ($\beta=0.5,\lambda=0$, top-left panel), flatter profile ($\beta=4,\lambda=0$, top-right panel), and off-centered profile ($\beta=1,\lambda=4$, bottom panels). For the off-centered profile, the early evolution is shown in the bottom-left panel, while the late evolution is shown in the bottom-right panel. In all cases, the perturbation amplitudes are chosen such that $\delta_{m}-\delta_{c}\approx 0.00226$.
Second, we consider a smoothed top-hat profile that transitions more sharply from $\bar{K}=1$ to $\bar{K}=0$:
$K(r)=\mathcal{A}\times\left\{\begin{aligned} &1\quad&\text{if}\quad r\leq\Delta_{1}\\ &\left(1+\frac{(r-\Delta_{1})^{2}}{2\Delta_{2}^{2}}\right)\exp\left(-\frac{(r-\Delta_{1})^{2}}{2\Delta_{2}^{2}}\right)\quad&\text{if}\quad r>\Delta_{1}\\ \end{aligned}\right$., (60)
and its first derivative is
$K^{\prime}(r)=\mathcal{A}\times\left\{\begin{aligned} &0\quad&\text{if}\quad r\leq\Delta_{1}\\ &-\frac{(r-\Delta_{1})^{3}}{2\Delta_{2}^{4}}\exp\left(-\frac{(r-\Delta_{1})^{2}}{2\Delta_{2}^{2}}\right)\quad&\text{if}\quad r>\Delta_{1}\\ \end{aligned}\right.$. (61)
Here, $\Delta_{1}$ determines the size of the top-hat region, while $\Delta_{2}$ controls the width of the transition region. We vary the value of $\Delta_{1}$ as a free parameter, while $\Delta_{2}$ is related to $\Delta_{1}$ and $r_{i}$ by
$\Delta_{2}=\sqrt{-\frac{\Delta_{1}^{2}}{4}+\frac{\Delta_{1}r_{i}}{2}-\frac{r_{i}^{2}}{4}+\frac{\sqrt{\Delta_{1}^{4}-8\Delta_{1}^{3}r_{i}+18\Delta_{1}^{2}r_{i}^{2}-16\Delta_{1}r_{i}^{3}+5r_{i}^{4}}}{4}}.$ (62)
We consider two cases: $\Delta_{1}=4R_{H}$ and $\Delta_{1}=8R_{H}$. The curvature perturbation profiles for both cases are shown as dashed lines in Fig. 11. The corresponding formation thresholds are $\delta_{c}\approx 0.53588\pm 1\times 10^{-5}$ and $\delta_{c}\approx 0.58689\pm 1\times 10^{-5}$, respectively. We choose the perturbation amplitudes to satisfy $\delta_{m}-\delta_{c}\approx 0.00226$, and Fig. 13 displays the time evolution of the energy density contrast for these two cases. The left column shows the early evolution, while the right column shows the late evolution. During the early evolution, the top-hat region gradually shrinks as matter flows inwards, increasing the central overdensity. The PBH forms at $t\approx 22.3t_{m}$ and $t\approx 18.5t_{m}$, respectively. After PBH formation, a bounce occurs, leading to the formation of a compression wave similar to that observed in the near-critical case for the Gaussian profile. Comparing the two cases reveals that $\Delta_{1}$ has little effect on compression wave formation, as both cases exhibit similar behavior.
Figure 13: Time evolution of the energy density contrast for smoothed top-hat profiles with $\Delta_{1}=4R_{H}$ (top row) and $\Delta_{1}=8R_{H}$ (bottom row). The early evolution is shown in the left column, while the late evolution is shown in the right column. In all cases, the perturbation amplitudes are chosen such that $\delta_{m}-\delta_{c}\approx 0.00226$.
Finally, we measure the thickness and energy of the sound shells, as described in subsection III.3 and III.4, for the different curvature perturbation profiles considered above. The measurement time is $t=200t_{m}$, after which the shell thickness and energy remain nearly constant. The results are summarized in Table LABEL:tab:summary. Our findings suggest that the main factor influencing the properties of the sound shell is the steepness of the perturbation profiles, while the influence of offset or top-hat regions is relatively minor. Steeper profiles tend to produce thicker and more energetic sound shells.
Profile $\delta_{c}$ $d_{+}/R_{H}$ $d_{-}/R_{H}$ $E_{+}/M_{H}$ $E_{-}/M_{H}$
Gaussian